if 4l^2-5m^2+6l+1=0 then the line lx+my+1=0 touches a fixed circle of centre (3,0) and radius


a)3^1/2 b)2^1/2 c)5^1/2 d)7^1/2

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Vaibhav

4l2-5m2+6l+1=0

 5( l2 +m2) =9l2 +6l +1

or 5( l2 +m2) =(3l+1)2

 or v5v( l2 +m2) =|(3l+1)|

now length of perpendicular from(3,0)  on line lx +my +1 =0

         r = |3l+1|/v( l2 +m2)

            from above rresult

    r = v5

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin